To determine the number of deer in a game preserve, a conservationist catches 788 deer, tags them and lets them loose. Later, 348 deer are caught, 174 of them are tagged. How many deer are in the preserve?
_________ deer are in the preserve?

Answer by checkley71(8403)

174/348=788/x
174x=274,224
x=274,224/174
x=1,576 is the estimated size of the herd.